A mall supervisor has found that mall sellers, on the average, complete a certain task in 10 minutes. If the times required to complete the task are approximately normally distributed with a standard deviation of 3 minutes, find: (3 marks each)
a. The proportion of mall sellers completing the task in less than 4 minutes
b. The proportion of mall sellers requiring more than 5 minutes to complete the task
c. The probability that a mall seller who has just been assigned the task will complete it within 3 minutes
1
Expert's answer
2020-09-14T18:53:30-0400

Given that the tasks are approximately normally distributed,


Population mean "\mu" = 10

Standard deviation "\sigma" = 3

x = time taken by the mall supervisors to complete a task


The formula for standard normal variable Z is given by,

Z = "\frac {x-\mu}{\sigma}"


Answer a)


To find the proportion of mall sellers completing the task in less than 4 minutes is given by,


p(x<4) = "p(\frac {x-\mu}{\sigma}<\frac {4-\mu}{\sigma})=p(Z<\frac {4-10}{3})=p(Z<-2)"


From the standard normal distribution table we get the probability for p(x<4) = 0.0228


Answer b)


To find the proportion of mall sellers requiring more than 5 minutes to complete the task


p(x>5) = "p(\frac {x-\mu}{\sigma}>\frac {5-\mu}{\sigma})=p(Z>\frac {5-10}{3})=p(Z>-1.6666)"


From the standard normal distribution table we get the probability for p(x>5) = 1 - 0.0485 = 0.9522


Answer b)


To find the probability that a mall seller who has just been assigned the task will complete it within 3 minutes


p(x<3) = "p(\frac {x-\mu}{\sigma}<\frac {3-\mu}{\sigma})=p(Z<\frac {3-10}{3})=p(Z<-2.333)"


From the standard normal distribution table we get the probability for p(x<3) = 0.0099